OpenStudy (anonymous):

3. The sum of three consecutive even integers is 84. What is the least of these numbers? (Points : 1) -34 -30 -26 30

OpenStudy (anonymous):

let x be the least of these numbers... since they're consecutive, the next one would be (x+2), and the next one after that is (x+4)... since their sum is 84, you can write it out as: x + (x+2) + (x+4) = 84 now all u gotta do is solve....

OpenStudy (anonymous):

How do I solve that???

OpenStudy (anonymous):

combine like terms and isolate x

OpenStudy (anonymous):

I dont know how to do that

OpenStudy (anonymous):

x + x + x are all the same.. so add them up = 3x... same with the 4 2

OpenStudy (anonymous):

now you should have 3x + 6 = 84. now you want your x term on one side of the equation and the numbers on the other side.

OpenStudy (anonymous):

is it 18?

OpenStudy (anonymous):

subtract 6 from both sides... thus 3x + 6 - 6 = 84 - 6

OpenStudy (anonymous):

so 3x = 78. now you need to get rid of that 3.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

you can divide by the coefficient of x.. in this case 3. thus 3x/3 = 78/3

OpenStudy (anonymous):

that will give you x = the result of 78/3

OpenStudy (anonymous):

26? :)

OpenStudy (anonymous):

oh wait -26?

OpenStudy (anonymous):

yep.. now check.. 26, + 28 + 30?

OpenStudy (anonymous):

i dont think someone can be negative years old.

OpenStudy (anonymous):

positive 26 is NOT an option :(

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

ooh i thought it was an age question. sorry

OpenStudy (anonymous):

its ok :/

OpenStudy (anonymous):

all those negatives dont make any sense...it is .. i think.. impossible for you to take a negative number with 3 consecutive even integers and add them up to 24.

OpenStudy (anonymous):

and 30 doesnt work

OpenStudy (anonymous):

Nope it was 30

OpenStudy (anonymous):

I just got the answers

OpenStudy (anonymous):

30+32+34 is 96?

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

30 + 32 + 34 \(\large \ne 84\) so the answer cannot be 30.... the answer to your question is 26... 26 + 28 + 30 = 84 looks like the correct choice is not there or the problem was mistyped... this is your orginal question: 3. The sum of three consecutive even integers is 84. What is the least of these numbers? (Points : 1) -34 -30 -26 30